OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (anonymous):

no u

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

pleeeeeaseeeeeeeeeee? I am having to retake all this :p And i dont remember a thing!! Its from the begining of the year :P

OpenStudy (anonymous):

x+y=170 x-y=26 __________ So x=26+y 26+y+y=170 solve for y, plug and chug the first eqn to get x

OpenStudy (anonymous):

wat???

OpenStudy (anonymous):

woops messed up.

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

nvm no I didn't

OpenStudy (anonymous):

You know that x and y together are 170 You know the difference is 26 So you know x or y being added to 26 is y or x respectively (doesnt matter) So you sub that into your first eqn. Then solve.

OpenStudy (anonymous):

\[x+y=170\] \[x-y=26\] add to get \[2x=196\] \[x=148\]

OpenStudy (anonymous):

x+y=170 and y-x=26 y-x=26 y=26+x x+26+x=170 2x=170-26 2x=144 x=72 x+y=170 72+y=170 y=170-72 y=98

OpenStudy (anonymous):

that is the correct answer:)

OpenStudy (anonymous):

and since \[x=148\] \[148+y=170\] \[y=170-148=22\]

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

\[\frac{196}{2}=98\]you are correct and i am wrong.

OpenStudy (anonymous):

thank you satellite73. it is okay.

OpenStudy (anonymous):

sometimes i have a hard time dividing by 2!

OpenStudy (anonymous):

lol, that is why, being careful will help us arrive to the answer correctly.